OpenStudy (anonymous):

What is the solution set for |2x – 8| ≥ 10?

OpenStudy (anonymous):

x >=8 x <=-1

OpenStudy (anonymous):

|2x – 8| ≥ 10 2x + 8 ≥ 10 -8 -8 2x ≥ 2 /2 /2 x ≥ 1 |2x – 8| ≥ 10 2x + 8 ≥ -10 - 8 -8 2x ≥ - 18 /2 /2 x ≥ -9

OpenStudy (chriss):

\[(\infty,-1]\cup[9,\infty)\]

OpenStudy (anonymous):

ss={x/x>or = 9}

OpenStudy (anonymous):

But these were the choices given A. x ≤ –1 or x ≥ 9 B. x ≤ –1 or x ≤ 9 C. x ≥ 9 or x ≥ –9 D. x ≤ 1 or x ≥ 9

OpenStudy (anonymous):

ChrisS has the right answer. He wrote it in interval notation. The answer he gave you means for all the values of x which are less than or equal to -1 or all the values or x which are greater than or equal to 9. you should be able to translate those sentences. The other two are wrong. Mitosuki forgot to flip the inequality when setting the problem up with -10. The other guy is just wrong.

OpenStudy (anonymous):

p.s. you should give ChrisS a medal